Ahmad finished his English homework in 1/5 hours. then he completed his math assignment in 3/4 hours. how much more time did Ahmad spend on his math assignment?


1
Expert's answer
2020-01-23T09:39:48-0500

We need to substract "\frac{1}{5}" from "\frac{3}{4}" using the Common Denominator Method to make the denominators the same.


"\frac{3}{4}-\frac{1}{5}=\frac{3\cdot{5}}{4\cdot{5}}-\frac{1\cdot{4}}{5\cdot{4}}=\frac{15}{20}-\frac{4}{20}=\frac{11}{20}"



Answer: Ahmad spent "\frac{11}{20}" hours more on his math assignment.